OpenStudy (anonymous):

X/5-g=a Solve for x. Please explain cause I seriously don't get it. Thx! It's x over 5

OpenStudy (anonymous):

is it x over 5 or x divided by 5?

OpenStudy (anonymous):

Over

OpenStudy (anonymous):

if x/5 - g = a, then if I am correct x should equal x= a*5 + g, because if x divided by 5 = a -g then a times 5 plus g should equal x. D o you understand?